# Online Multivariate Changepoint Detection: Leveraging Links With Computational Geometry

### v2

@@ 3.1 Mathematical Formulation
Functional pruning ideas give the potential of an efficient approach to recursively calculate the GLR statistic for a change at time $n+1$ given calculations at time $n$ [34]. The idea is to restrict the set of possible changepoints prior to $n+1$ that we need to consider. Working out which changepoint locations we need to consider is obtained by viewing the GLR statistic as a function of the post-changepoint parameter difference $\mu=\eta_{2}-\eta_{1}$ (see Equation (8)), and finding the set of changepoint locations which contribute to this, i.e. for which the GLR statistic is maximum for some value of $\mu$.
In more details, at step $n$ of a functional pruning algorithm, we consider a list of possible changepoints $\tau$ of the data from $x_{1}$, $x_{2}$to $x_{n}$. Each changepoint is represented by a function $f_{\tau}$ of the last segment parameter $\mu$ and a set of $\mu$ values for which it is maximising the likelihood: $Set_{\tau}=\{\mu|\forall\tau^{\prime}\neq\tau,\ f_{\tau}(\mu)>f_{\tau^{\prime}}(\mu)\}$. To proceed to the next data point $n+1$ the update is informally done as follows.
@@
    We include in the list of segmentations the one with a last change at $n$ which is represented by a function $f_{n}$.
@@
    We compare all other functions to this new function and accordingly restrict the set of $\mu$ for which they are optimal. That is we compute the intersection $Set_{\tau}\leftarrow Set_{\tau}\cap\{\mu|f_{\tau}(\mu)>f_{n}(\mu)\}$ for all $\tau<n,$ and the intersection $Set_{n}\leftarrow\cap_{\tau}\{\mu|f_{n}(\mu)>f_{\tau}(\mu)\}.$
@@
The left column of Figure 1 provides a graphical representation of these steps. The pruning of the third step is valid because whatever the data points after $n+1$ might be these functions (or equivalently changepoints) can never maximise the log-likelihood. Formally, we now consider a set $\mathcal{T}$ indexing functions $f_{\tau}$ such that
$f_{\tau}(\mu)=a_{\tau}A(\mu)-2\langle\mu,b_{\tau}\rangle+c_{\tau}\,,$ (9)
where $a_{\tau}$, $c_{\tau}$ are in $\mathbb{R}$, and $b_{\tau}$ is in $\mathbb{R}^{p}$ and $A$ is a convex function from $\mathcal{D}_{A}$, a subset of $\mathbb{R}^{p}$. We define $Im(A)$ as the image of the function $A$. The idea is that at the current step of a functional pruning algorithm we are storing functions of the post-changepoint parameter difference, here denoted by $\mu$. Function $f_{\tau}$ corresponds to the log-likelihood for a segmentation with a change at $\tau$ and with parameter difference $\mu$. This function will depend on $\mu$ through the log-likelihood contribution of the data in this segment, which will impact all the coefficients and the maximum log-likelihood for data in earlier segments. The functional form of $f_{\tau}$ will depend on the specific likelihood model. As an example, for detecting a change in the mean of Gaussian data with $\eta_{1}=0$, we would have $A(\mu)=||\mu||^{2}$, with $Im(A)=\mathbb{R}^{+}$ and $\mathcal{D}_{A}=\mathbb{R}^{p}$.
We then define
@@
and, as informally stated, our goal is to recover all the indices $\tau$ maximising $F_{\mathcal{T}}$ for at least one value of $\mu$. Formally, we define:
@@
Remark 2.
Index set $\mathcal{F}_{\mathcal{T}}$ is defined using a strict inequality. As function $A$ is strictly convex, (and assuming that for all $\tau\neq\tau^{\prime}$ we cannot have simultaneously $a_{\tau}=a_{\tau^{\prime}}$, $b_{\tau}=b_{\tau^{\prime}}$ and $c_{\tau}=c_{\tau^{\prime}}$) equality can be obtained only over a set of zero measure, consequently this case can be ignored.
@@ 3.2 Linearisation of the Functional Pruning Problem
$g_{\tau}(\lambda,\mu)=a_{\tau}\lambda-2\langle\mu,b_{\tau}\rangle+c_{\tau}.$
@@
Without any further assumptions on the function $A$, the set $\mathcal{G}_{\mathcal{T}}$ could be much larger than $\mathcal{F}_{\mathcal{T}}$. The following theorem demonstrates that equality holds for strictly convex power functions.
@@
If for all $\tau\in\mathcal{T}$, $c_{\tau}=0$ and $A(\mu)=\|\mu\|^{q}$ with $q>1$, then $\mathcal{F}_{\mathcal{T}}=\mathcal{G}_{\mathcal{T}}$.
The proof of this Theorem can be found in Supplementary Material A.1.
Remark 3.
In particular, with $q=2$ we get the Gaussian loss function with $\eta_{1}=0$. Assuming $\eta_{1}=0$ is not restrictive because if $\eta_{1}$ is not 00 we can consider the translated data $(x_{t}-\eta_{1}).$
Remark 4.
A more general approach with milder conditions for function $A$ is proposed in Supplementary Material B. One sufficient condition consists of the strong convexity of the function $A$ with $A(0)=0$.
@@ 3.3 Half-space intersections and functional pruning
$H_{\tau}=\{\kappa,\lambda,\mu\ |\ h_{\tau}(\kappa,\lambda,\mu)=a_{\tau}\lambda-2\langle b_{\tau},\mu\rangle+c_{\tau}-\kappa\leq 0\}.$
@@
    the hyperplane defined by $h_{\tau}(\kappa,\lambda,\mu)=0$ is a supporting hyperplane of the set $\cap_{\tau}H_{\tau}$.
@@
The set $\cap_{\tau}H_{\tau}$ defines an unbounded polyhedron as taking any $\kappa\geq\max_{\tau}\{c_{\tau}\}$, $\lambda=0$, and $\mu=0$ we solve all inequalities.
In Figure 1 we illustrate our linearised functional problem for the special case where, for all $\tau$ in $\mathcal{T}$, $c_{\tau}=0$.
Computing the intersection of $|\mathcal{T}|$ half-spaces is a well-studied problem in computational geometry. It is closely related to the convex hull problem by projective duality [43] and several algorithms have been proposed (see for example [44, 45, 46, 47, 48, 49]). See [50] for a recent short introduction to this topic.
(a) Iteration $n=3$
(b) Iteration $n=4$
Figure 1: Example of the functional pruning problem and linearised functional problem over times: $f_{i}$ and $\operatorname{proj}\{\max_{i}\{f_{i}(\mu)\}\}$ functions (left); $g_{i}$ and $\operatorname{proj}\{\max_{i}\{g_{i}(\mu,\lambda)\}\}$ functions( middle); the level surfaces of $h_{i}$ functions, where each surface is for a fixed $\kappa$ value (right) for changepoint candidates $i\in\mathcal{T}$. We simulate time series $\{x_{t}\}_{t=1,2,..}\sim\mathcal{N}(0,1)$ and consider two time-steps, $n=3$ and $n=4$. We define the coefficients $a_{i}$ as $i$, $b_{i}$ as $\sum_{t=1}^{i}x_{t}$ and $c_{i}$ as zero. Note that the change $i=2$ associated with quadratics $f_{2}$ and surface $g_{2}$ is pruned at $n=4$.
3.4 Projective-duality
In this section we study the special case where for all $\tau$ in $\mathcal{T}$ $c_{\tau}=0.$ The following theorem states that in this case the half-space intersection problem in dimension $p+2$ is dual to a convex hull problem in dimension $p+1$.
@@ 3.4 Projective-duality
If for all $\tau$ in $\mathcal{T}$, $c_{\tau}=0$, then $\tau$ being in $\mathcal{G}_{\mathcal{T}}$ is equivalent to the point $(a_{\tau},b_{\tau})$ in $\mathbb{R}^{p+1}$ being a vertex of the convex hull of $\{(a_{\tau},b_{\tau})\}_{\tau\in\mathcal{T}}$.
The proof can be found in Supplementary Material A.2.
We illustrate the points on the convex hull of Theorem 4 for $p=1$ and $p=2$ in Figure 2. As will be seen in Section 4 this is relevant for the online detection of a single changepoint (see also Equation (8)). Importantly, computing the convex hull of $|\mathcal{T}|$ points is a well-studied problem in computational geometry. Over the years many algorithms have been developed [51, 52, 48, 53, 54, 55, 56]. In the rest of this paper we will only use the QuickHull algorithm of [48] and its implementation in the qhull library http://www.qhull.org/. The QuickHull algorithm works in any dimension and is reasonably fast for our application up to $p=5$.
(a) Dimension $p=1$
(b) Dimension $p=2$
Figure 2: Plot of $S(i)=\sum_{t=1}^{i}x_{t}$ as a function of $i$ (dark blue) and the convex hull of $\left\{\left(a_{i}=i,b_{i}=S(i)\right)\right\}_{i\in\{1,\dots,n-1\}}$ (dashed blue line) for dimensions $p=1$ and $2$ The vertices of the convex hull $\tau\in\mathcal{T}$ are marked in red. We simulated time series $\{x_{t}\}_{t\in\{1,\dots,n\}}$ with $n=51$ data points (without changepoints, $x_{t}\sim\mathcal{N}_{p}(0,I_{p})$ i.i.d.).
4 On The Detection of a Single Changepoint Online
4.1 From Functional Description to Convex Hull
$P(\tau)=\left(\tau,\sum_{t=1}^{\tau}x_{t}\right).$
The following theorem, based on Theorem 4 essentially states that the changepoint maximising the likelihood (6) for some value $\eta_{1}$ and $\eta_{2}$ corresponds to a points on a particular convex hull.
Theorem 5.
Assume the log-likelihood, $\ell_{\tau,n}(\eta_{1},\eta_{2})$ of a change at $\tau$ in $\{1,\ldots,n-1\}$ can be written as in Equation (6). Then, for any $\eta_{1}$ and $s\leq p$ taking $\hat{\tau}=\hat{\tau}^{s}(\eta_{1})$ or $\hat{\tau}^{s}(.)$ we have that $P(\hat{\tau})$ is on the convex hull of $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$.
The proof of this theorem can be found in Supplementary Material A.3.
Theorem 5 is coherent with the notion of pruning as in the FOCuS algorithm [34]. If $P(\tau_{0})$ is strictly inside the hull (i.e. not on the boundary of the hull $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$) at time $n$, then it is certainly inside the hull at any future time $n^{\prime}\geq n$ (i.e. not on the boundary of the hull $\{P(\tau)\}_{\tau\in\{1,\ldots,n^{\prime}-1\}}$). Thus based on Theorem 5 the likelihood associated with the change at time $\tau_{0}$ will never be optimal after time $n$.
From a computational perspective, Theorem 5 is useful only if the set of points on the convex hull is significantly smaller than $n$. In the following subsections, we will prove using [57] that, under natural assumptions, the number of vertices of the convex hull of $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$ is of order $2(\log n)^{p}$. This shows that for sufficiently large $n$ the number of candidate changepoints one needs to store to optimise the likelihood (or the likelihood ratio test) is much smaller than $n$.
4.2 Bound on The Number of Changepoint Candidates
In this subsection using the recent paper [57], we provide the expected number of faces and vertices of the convex hull $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$, assuming that the components of $x_{\tau}$, $x_{\tau}^{d}$ (with $d=1,\dots,p$), are independent and identically distributed with a continuous distribution. In turn, using Theorem 4 this gives us an upper bound on the expected number of changepoints that can maximise the likelihood function (6) for some value of $\eta_{1}$ and $\eta_{2}$. From a computational perspective, this bound will be useful to derive guarantees on the computational complexity of the procedure (see Algorithms 1 and 3).
To apply the result of [57] we need to prove that the random walk $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}=\{(\tau,\sum_{t=1}^{\tau}x_{t})\}_{\tau\in\{1,\ldots,n-1\}}$ satisfies two properties, $(Ex)$ and $(GP)$ defined below.
Lemma 1.
Assuming that all $x_{\tau}^{d}$ are independent and identically distributed with a continuous distribution, the properties $(Ex)$ and $(GP)$ of [57] are satisfied.
The proof of this lemma can be found in Supplementary Material A.4.
We can now state the expected number of faces and vertices of the convex hull of $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$.
Theorem 6.
Assuming all $x_{\tau}^{d}$ (with $d=1,\dots,p$) are independent and identically distributed with a continuous distribution, then the expected number of faces (denoted as $U_{n}^{p}$) and vertices (denoted as $V_{n}^{p}$) of the convex hull of $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$ are the following:
@@ 4.2 Bound on The Number of Changepoint Candidates
where $\begin{bmatrix}n\\ m\\ \end{bmatrix}$ are the Stirling numbers of the first kind [58].
Proof.
Based on Lemma 1 properties $(Ex)$ and $(GP)$ hold and we can apply formula (2) from [57] on the expected values of $U_{n}^{p}$ and $V_{n}^{p}.$
@@
The proof of Corollary 1 follows from Remark 1.4 of [57]. The theorem and corollary provide an expected upper limit on the number of points, or candidate locations for changepoints, that need to be evaluated to compute the CUSUM likelihood ratio test (6) exactly without resorting to approximations. Although this limit increases as $\mathcal{O}(\log(n)^{p})$ and thus as a power of $p$, it is not overly restrictive. In some online applications, we often encounter situations where $n>>p$, making an exact implementation feasible. In the following paragraph, we will evaluate scenarios up to $5$ dimensions.
Remark 6.
Theorem 6 gives the expectation for i.i.d. observations. One might wonder what happens if there is a changepoint in the distribution at $i<n.$ Note that the points on the hull $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$ are necessarily on the hull of $\{P(\tau)\}_{\tau\in\{1,\ldots,i\}}$ or on the hull of $\{P(\tau)\}_{\tau\in\{i+1,\ldots,n-1\}}.$ Thus applying the theorem on all points from $1$ to $i$ and from $i+1$ to $n-1$ we expect at most $\mathbb{E}(V_{i+1}^{p})+\mathbb{E}(V_{n-i}^{p})\leq 2\mathbb{E}(V_{n}^{p})$ points on the hull (as the expectation of $V_{i}^{p}$ is increasing with $i$). Thus if there is a changepoint (and assuming we haven’t stop because the GLR statistics is lower than the specified threshold) the expected number of candidate changepoints is still in $\mathcal{O}(\log(n)^{p}).$
Empirical validation of the bound
We empirically tested the validity of Theorem 6. We use the Gaussian distribution as an example of a distribution with continuous density and independent dimensions. To be specific, for $p$ from $1$ to $5$, we simulated time series $x_{1:n}$, with $x_{t}\sim\mathcal{N}_{p}(0,I_{p})$ i.i.d. We varied the time series length from $2^{10}+1(\approx 10^{3})$ to $2^{23}+1(\approx 8\times 10^{6})$ and simulated $100$ data sets for each length. We used the QuickHull algorithm [48] implemented in the R package geometry [59] to evaluate the number of faces and vertices of the convex hull of $\{P((\tau)\}_{\tau\in\{1,\ldots,n-1\}}$. It can be seen in Figure 4 that the observed number of faces and vertices are close to their theoretical expectations (from Theorem 6). Additional details on the calculations of the Stirling numbers are provided in Supplementary Material C.
The proof of Theorem 6 depends on the fact that the distribution is continuous. We thus empirically tested the robustness of the expectations to this assumption by considering time series drawn with a discrete i.i.d. Poisson distribution. The results closely resemble the Gaussian case and details can be found in Supplementary Material E.2.
4.3 MdFOCuS(0): a Simple Online Algorithm

### v3

@@ 3.1 Mathematical Formulation
Functional pruning ideas give the potential of an efficient approach to recursively calculate the GLR statistic for a change at time $n+1$ given calculations at time $n$ Romano et al. (2023). The idea is to restrict the set of possible changepoints prior to $n+1$ that we need to consider. Working out which changepoint locations we need to consider is obtained by viewing the GLR statistic as a function of the change parameter $\mu=\eta_{2}-\eta_{1}$, and finding the set of changepoint locations which contribute to this, i.e. for which the GLR statistic is maximum for some value of $\mu$.
In more detail, at step $n$ of a functional pruning algorithm, we consider a list of possible changepoints $\tau$ of the data from $x_{1}$, $x_{2}$to $x_{n}$. Each changepoint is represented by a function $f_{\tau}$ of the last segment parameter $\mu$ and a set of $\mu$ values for which it is maximising the likelihood: $Set_{\tau}=\{\mu\;|\;\forall\ \tau^{\prime}\neq\tau,\ f_{\tau}(\mu)>f_{\tau^{\prime}}(\mu)\}$. To proceed to the next data point $n+1$ the update is informally done as follows.
@@
    We add $n$ to the list of changepoints, with this represented by a function $f_{n}$.
@@
    We compare all other functions to this new function and accordingly restrict the set of $\mu$ for which they are optimal.
@@
The left column of Figure 1 provides a graphical representation of these steps. The pruning of the third step is valid because, whatever the data points after $n+1$ might be, these functions (or equivalently changepoints) can never maximise the log-likelihood. Formally, define functions $f_{\tau}$, for $\tau$ in some set $\mathcal{T}$, such that
$f_{\tau}(\mu)=a_{\tau}A(\mu)-2\langle b_{\tau},\mu\rangle,$ (8)
where $a_{\tau}$ is in $\mathbb{R}$, and $b_{\tau}$ is in $\mathbb{R}^{p}$ and $A$ is a convex function from $\mathcal{D}_{A}$, a subset of $\mathbb{R}^{p}$. We define $Im(A)$ as the image of the function $A$. Function $f_{\tau}$ corresponds to the log-likelihood for a segmentation with a change at $\tau$ and with parameter $\mu$. This function will depend on $\mu$ through the log-likelihood contribution of the data in the segment after $\tau$, with this data defining the co-efficients $a_{\tau}$ and $b_{\tau}$. We then define
@@
and our goal is to recover all the indices $\tau$ maximising $F_{\mathcal{T}}$ for at least one value of $\mu$. Define:
@@
This is defined using a strict inequality. As function $A$ is strictly convex, (and assuming that for all $\tau\neq\tau^{\prime}$ we cannot have simultaneously $a_{\tau}=a_{\tau^{\prime}}$ and $b_{\tau}=b_{\tau^{\prime}}$) equality can be obtained only over a set of zero measure, and can be ignored.
@@ 3.2 Linearisation of the Functional Pruning Problem
$g_{\tau}(\lambda,\mu)=a_{\tau}\lambda-2\langle b_{\tau},\mu\rangle.$
@@
Without any further assumptions on the function $A$, the set $\mathcal{G}_{\mathcal{T}}$ could be much larger than $\mathcal{F}_{\mathcal{T}}$. The following theorem demonstrates that equality holds for an important class of functions $A(\mu)$.
@@
If $A(\mu)=\|\mu\|^{q}$ with $q>1$, then $\mathcal{F}_{\mathcal{T}}=\mathcal{G}_{\mathcal{T}}$.
In particular, with $q=2$ we get the Gaussian loss function with $\eta_{1}=0$. Assuming $\eta_{1}=0$ is not restrictive because if $\eta_{1}$ is not 00 we can consider the translated data $(x_{t}-\eta_{1})$. A more general result, with milder conditions for function $A$, is proposed in Supplementary Material D. One sufficient condition consists of the strong convexity of the function $A$ with $A(0)=0$.
@@ 3.3 Half-space intersections and functional pruning
$H_{\tau}=\{\kappa,\lambda,\mu\ |\ h_{\tau}(\kappa,\lambda,\mu)=a_{\tau}\lambda-2\langle b_{\tau},\mu\rangle-\kappa\leq 0\}.$
@@
    $h_{\tau}(\kappa,\lambda,\mu)=0$ defines a supporting hyperplane of the set $\cap_{\tau}H_{\tau}$.
The set $\cap_{\tau}H_{\tau}$ defines an unbounded polyhedron as taking any $\kappa\geq\max_{\tau}\{c_{\tau}\}$, $\lambda=0$, and $\mu=0$ we solve all inequalities. In Figure 1 we illustrate our linearised functional problem.
Computing the intersection of $|\mathcal{T}|$ half-spaces is a well-studied problem in computational geometry. It is closely related to the convex hull problem by projective duality Avis and Bremner (1995), and several algorithms have been proposed (see for example Preparata and Muller, 1979; Brondsted, 1983; Barber et al., 1996; O’Rourke, 1998). See Seidel (2017) for a recent short introduction to this topic.
Figure 1: Example of the functional pruning problem and linearised functional problem over times: $f_{i}$ and $\operatorname{proj}\{\max_{i}\{f_{i}(\mu)\}\}$ functions (left); $g_{i}$ and $\operatorname{proj}\{\max_{i}\{g_{i}(\mu,\lambda)\}\}$ functions (middle); the level surfaces of $h_{i}$ functions, where each surface is for a fixed $\kappa$ value (right) for changepoint candidates $i\in\mathcal{T}$. We simulate time series $\{x_{t}\}_{t=1,2,..}\sim\mathcal{N}(0,1)$ and consider two time-steps, $n=3$ (top two rows) and $n=4$ (bottom two rows). We define the coefficients $a_{i}$ as $i$, $b_{i}$ as $\sum_{t=1}^{i}x_{t}$. The change $i=2$ associated with quadratics $f_{2}$ and surface $g_{2}$ is pruned at $n=4$.
3.4 Projective-duality
The following theorem states that in our case the half-space intersection problem in dimension $p+2$ is dual to a convex hull problem in dimension $p+1$.
Theorem 4.
An index $\tau$ being in $\mathcal{G}_{\mathcal{T}}$ is equivalent to the point $(a_{\tau},b_{\tau})$ in $\mathbb{R}^{p+1}$ being a vertex of the convex hull of $\{(a_{\tau},b_{\tau})\}_{\tau\in\mathcal{T}}$.
We illustrate the points on the convex hull of Theorem 4 for $p=1$ and $p=2$ in Figure 2. As will be seen in Section 4, this is relevant for the online detection of a single changepoint, see (7). Importantly, computing the convex hull of $|\mathcal{T}|$ points is a well-studied problem in computational geometry. Over the years many algorithms have been developed Chand and Kapur (1970); Preparata and Hong (1977); Miller and Stout (1988); Barber et al. (1996); Kenwright (2023). In the rest of this paper, we will use the QuickHull algorithm of Barber et al. (1996) and its implementation in the qhull library http://www.qhull.org/. The QuickHull algorithm works in any dimension and is reasonably fast for our application up to $p=5$.
Figure 2: Plot of $S(i)=\sum_{t=1}^{i}x_{t}$ as a function of $i$ (dark blue) and the convex hull of $\left\{\left(a_{i}=i,b_{i}=S(i)\right)\right\}_{i\in\{1,\dots,n-1\}}$ (dashed blue line) for dimensions $p=1$ (left) and $2$ (right). The vertices of the convex hull $\tau\in\mathcal{T}$ are marked in red. We simulated time series $\{x_{t}\}_{t\in\{1,\dots,n\}}$ with $n=51$ data points with $x_{t}\sim\mathcal{N}_{p}(0,I_{p})$.
4 On The Detection of a Single Changepoint Online
4.1 From the Functional Description to a Convex Hull
$P(\tau)=\left(\tau,\sum_{t=1}^{\tau}x_{t}\right).$ (10)
The following theorem, based on Theorem 4 essentially states that the changepoint maximising the likelihood (5) for some value $\eta_{1}$ and $\eta_{2}$ corresponds to a point on a particular convex hull.
Theorem 5.
Assume the log-likelihood, $\ell_{\tau,n}(\eta_{1},\eta_{2})$ of a change at $\tau$ in $\{1,\ldots,n-1\}$ can be written as in (5). Then, for any $\eta_{1}$ and $s\leq p$ taking $\hat{\tau}=\hat{\tau}^{s}(\eta_{1})$ or $\hat{\tau}^{s}(.)$ we have that $P(\hat{\tau})$ is on the convex hull of $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$.
@@ 4.1 From the Functional Description to a Convex Hull
This result also extends to other test statistics that maximise the sparse log-likelihood for some $s$. This includes statistics that threshold the contribution from each component of the time series (see Section I.1 in Supplementary Material).
Theorem 5 is coherent with the notion of pruning as in the FOCuS algorithm Romano et al. (2023). If $P(\tau_{0})$ is strictly inside the hull, i.e. not on the boundary of the hull $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$ at time $n$, then it is inside the hull at any future time $n^{\prime}\geq n$. Thus, based on Theorem 5, the likelihood associated with the change at time $\tau_{0}$ will never be optimal after time $n$.
4.2 Bound on The Number of Changepoint Candidates
From a computational perspective, Theorem 5 is useful only if the set of points on the convex hull is significantly smaller than $n$. We now use recent results Kabluchko et al. (2017) to bound the expected number of faces and vertices of the convex hull $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$, assuming that the components of $x_{\tau}$, $x_{\tau}^{d}$ (with $d=1,\dots,p$), are i.i.d. with a continuous distribution. In turn, using Theorem 4, this gives us an upper bound on the expected number of changepoints that can maximise the likelihood function (5) for some value of $\eta_{1}$ and $\eta_{2}$. This bound will be useful to derive guarantees on the computational complexity of the procedure (see Algorithms 1 and 3).
@@ 4.2 Bound on The Number of Changepoint Candidates
Assuming all $x_{\tau}^{d}$ (with $d=1,\dots,p$) are i.i.d. with a continuous distribution, then the expected number of faces (denoted as $U_{n}^{p}$) and vertices (denoted as $V_{n}^{p}$) of the convex hull of $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}$ are the following:
@@
where $\begin{bmatrix}n\\ m\\ \end{bmatrix}$ are the Stirling numbers of the first kind Adamchik (1997).
@@
The proof of Corollary 1 follows from Remark 1.4 of Kabluchko et al. (2017). The theorem and corollary provide an expected upper limit on the number of points, or candidate locations for changepoints, that need to be evaluated to compute the CUSUM likelihood ratio test (5) exactly without resorting to approximations. This limit increases as $\mathcal{O}(\log(n)^{p})$ and so will give small bounds on the number of candidate changepoints for small $p$. We empirically illustrate the validity of Theorem 6 for i.i.d. Gaussian and Poisson data in supplementary material B.5 and F.1.
4.3 MdFOCuS: a Simple Online Algorithm